ADNAN MUHAMMED

Grade 12,

two electrostatic force between 200 microcoulomb and 500 microcoulomb placed in free space is 5 cm find the distance between two charges

two electrostatic force between 200 microcoulomb and 500 microcoulomb placed in free space is 5 cm find the distance between two charges

Grade:12

1 Answers

Rajat
5 years ago
Your question is incomplete, please provide the value of force.
Let force be F then we can use coulomb's law, 
F= kq1q2/r^2 
Here, 
Force= F
q1= 200×10-6C
q2= 500× 10-6C
r = .05 m
k = 9×109 Nm^2/C^2